Determine the focus and directrix of the parabola with the given equation. Sketch the

graph, indicate the focus, directrix, vertex, and axis of symmetry. (2 items, 20 points)

1. x^2 − 6x − 5y = −34

2. y^2 − 12x + 8y = −40


1
Expert's answer
2021-03-15T12:17:01-0400

Solution.

1) "x^2-6x-5y=-34."

"y=\frac{1}{5}(x-3)^2-5."

From here vertex of the parabola "(h;k)" will be "(3;5)."

"x=3" - axis of symmetry.

Focus "(h;k+\frac{1}{4a})," where "a=\frac{1}{5}." So, "(3;6.25)" is the focus of the parabola.

"y=k-\frac{1}{4a}" is the equation of directrix.

Therefore "y=3.75" is directrix.

2) "y^2-12x+8y=-40."

"(y+4)^2=12(x-2)."

From here "(2;-4)" is the vertex of parabola.

"y=-4" - axis of symmetry.

Focus "(h+p;k)," where "p=3." So, "(5;-4)" is the focus.

"x=-1" is the directrix of the parabola.